If the value you computed has an SI unit of 𝐶²m²/kgm², the quantity could be



a. potential difference c. electric field



b. dielectric constant d. capacitance

1
Expert's answer
2022-05-15T17:03:37-0400

SI units:

a) "\rm[V]=[kg\cdot m^2/(C\cdot s^2)]"

b) "\rm[\varepsilon]=[C^2\cdot s^2/(kg\cdot m^3)]"

c) "\rm[E]=[kg\cdot m/(C\cdot s^2)]"

d) "\rm[C]=[C^2\cdot s^2/(kg\cdot m^2)]"


Answer: d. capacitance